If 4.65 LL of CO2CO2 gas at 13 ∘C∘C at 777 mmHgmmHg is used, what is the final volume, in liters, of the gas at 39 ∘C∘C and a pressure of 735 mmHgmmHg, if the amount of CO2CO2 does not change?
Answer #1

Solution:-

Using equation,

​​​​​​(P1V1) / T1 = (P2V2) / T2

Where,

P1 = Initial Pressure

V1 = Initial Volume

T1 = Initial temperature

P2 = Final Pressure

V2 = Final Volume

T2 = Final temperature.

Given:- P1 = 777 mmHg ; V1= 4.65 L ; T1 = 13°C = (273.15 + 13) K = 286.15 K

P2 = 735 mmHg ; V2 = ? ; T2 = 39°C = (273.15 + 39) K = 312.15 K

Now,

(777 × 4.65) / (286.15) = (735 × V​​​​​​​2) / (312.15)

or, V2 = 5.362 L

Hence, final volume of CO2 = 5.362 L

----------------------------------------------